OpenStudy (anonymous):

Graph, using the slope and y intercept. y=9/4x+2

OpenStudy (katrinakaif):

Make a point at 2 on the y-intercept Then move 9 units up, and 4 units to the right, make another point. Keep moving 9 units up and 4 units right from the previous point, and make another point. Connect the dots =))

OpenStudy (anonymous):

there should be only two points right-to plot a line.

jimthompson5910 (jim_thompson5910):

Best way to do this is to plot two points and then connect them with a straight line. To get two points, plug in values of x to get corresponding values of y So when x = 0, y = 9/4(0)+2 = 0 + 2 = 2 This means that the point (0, 2) is on the line Also, when x = 4, y = 9/4(4)+2 =9+2 = 11 This tells us that (4, 11) is also on the line So plot the points (0, 2) and (4, 11), then connect them with a straight line to graph y = 9/4x+2